if sin^-1, x=2 pi/5 find the value of cos^-1
Answer by lwsshak3(11628)

if sin^-1, x=2 pi/5 find the value of cos^-1
sin^-1(x)=2π/5
cos^1(x)=π/2-2π/5=π/10
check:
sin(2π/5)≈0.951
cos(π/10)=0.951
note: sin(x)=cos(π/2-x)